# Generalized Pulse Design in Floquet Engineering: Application to Interacting Spin Systems

### v1

@@ I Introduction
Broadening the class of accessible quantum simulation models by improving the precision and applicability of engineering protocols will help experiments study models otherwise inaccessible to conventional numerical methods [1, 2, 3, 4, 5, 6, 7]. High frequency driving (Floquet engineering) with external fields has proven to be a useful experimental tool whereby otherwise static system Hamiltonians, $\hat{H}_{0}$, are perturbed with periodic external field pulses, $\hat{V}(t)$. The resulting driven Hamiltonian:
@@
and $T$ is a subcycle time defining periodicity such that $\hat{V}(t)$ is periodic in an integer multiple of $T$. $\omega^{-1}$ establishes a small expansion parameter compared to all other energy scales in $H_{0}$. Interesting properties of eigenstates of $\hat{H}_{\text{eff}}$ can then be observed experimentally.
Recent experiments with Rydberg atoms [8, 9, 10, 11], molecules [12], and trapped ions [13] have been able to dynamically engineer effective higher symmetry Heisenberg spin models from lower symmetry models. Ref. 9 used rectangular pulse profiles to engineer an effective $SU(2)$-symmetric Heisenberg interaction between a large number of Rydberg atoms initially lacking this symmetry in their interactions. Similarly, trapped ions were recently driven [13] from an Ising limit into a higher-symmetry Heisenberg regime to replicate the Haldane-Shastry model [14, 15]. Interest in engineering systems extends beyond transformations of underlying symmetry groups, or simulating $SU(2)$ Hamiltonians. For example, there have been proposals and experiments employing cold-atom systems to explore gauge theories and other field theories [16, 17, 18, 19], and trapped ions have been used to simulate complex open systems [20, 5].
The theory of Floquet engineering–and the associated high frequency expansions–has a long history (see Refs. [21, 22, 23, 24] for reviews). There are two common methods to perform controlled expansions in inverse frequency that allow for effective engineering of broad classes of Hamiltonians. The first is the high frequency stroboscopic Magnus expansion (SME) [25, 24]. SMEs are commonly used in derivations of $\hat{H}_{\text{eff}}$, particularly in the context of average Hamiltonian theory [26, 22]. SMEs can be pushed to high order in inverse frequency to derive non-trivial effective models, e.g., multi-spin interactions [27, 28], but often layer in additional assumptions. The second is the gauge invariant high frequency expansion (GIHFE) [29, 23, 24]. The GIHFE approach relies on the definition of a kick operator, $\hat{K}(t)$, from $\hat{V}(t)$ [23]. The effective Hamiltonian and new gauge transformed state follows from:
@@
From these expressions we see that derivation of $\hat{H}_{\text{eff}}$ and derivation of the new states $\phi(t)$ from the original $\psi(t)$ then requires a systematic inverse frequency expansion of $\hat{K}(t)$ [29, 23] in addition to the expansion for $\hat{H}_{\text{eff}}$.
GIHFEs, in contrast to SMEs, are often used to model sampling even at non-stroboscopic times by tracking the ensuing change in basis (gauge transformation) initiated by pulsing. Despite the distinction, one can recover average Hamiltonian theory from GIHFEs by restricting to stroboscopic sampling and applying the usual Magnus expansion [24]. The non-stroboscopic case can be constructed via a gauge transformation that depends on the initial sampling time. Assumptions underlying conventional GIHFE methods limit the space of available pulse profiles and therefore limit possible effective Hamiltonians.
We revisit the GIHFE formalism with the aim of broadening the set of available effective Hamiltonians in driven spin systems. The engineering method we construct generalizes the available pulse shapes considered in the existing literature [29, 23, 24]. We consider periodic pulse structures such that $\omega^{-1}$ defines the only small parameter allowing otherwise arbitrary pulse shapes. Our formalism then allows for large pulse strengths, any pulse width that respects the high frequency approximation, and non-stroboscopic sampling. The pulse shape freedom we find arises because we are able to leverage the commutation relations of the Pauli spin matrices to perform controlled series expansions approximating $\hat{K}(t)$. Hamiltonian engineering then reduces to a functional optimized over the entire space of pulse shapes so that the desired effective Hamiltonian is realized.
To exemplify our approach, we apply our formalism to decrease the symmetry implied by native spin-spin interactions, motivated by the desire to construct (weighted) graph states [30] in spin systems with native dipolar XXZ spin-spin interactions which characterize certain Rydberg atom [31, 32] and molecule [33, 34, 35, 36] interactions. Since XXZ interactions do not commute in general, the protocol required to build weighted graph states introduces complexities that the Ising-type interaction avoids [30, 37]. Requirements for building high precision weighted graph states with global pulses therefore motivates the specific example we use to demonstrate our formalism: engineering an XXZ interaction into an Ising interaction.
@@ II Spin Models
where $\hat{\sigma}^{(1)}_{j}$, $\hat{\sigma}^{(2)}_{j}$, and $\hat{\sigma}^{(3)}_{j}$ denote the usual $x$, $y$, and $z$ Pauli-spin matrices, respectively. $j$ and $j^{\prime}$ denote site (vertex) pairs in an arbitrary graph. $J^{1}_{jj^{\prime}}$, $J^{2}_{jj^{\prime}}$, and $J^{3}_{jj^{\prime}}$ denote spatially varying spin-spin interaction energies along the spin-$x$,$y$, and $z$ directions, respectively. In specifying example applications, we restrict our attention to XXZ-type models with:
@@
Equation (4) encompasses well-known spin models. For $J_{z}=0$, Eq. (4) reduces to an XY-model, denoted $H_{XY}$. For $J_{\perp}=0$, we retrieve the Ising model, denoted $H_{ZZ}$. Eq. (5) implies that we may write $\hat{H}_{0}^{\text{S}}$ as $\hat{H}_{XXZ}=\hat{H}_{XY}+\hat{H}_{ZZ}$. Other choices of the interaction constants give rise to additional models, such as the Heisenberg model ($J_{\perp}=J_{z}$). Our formalism will hold for general spin systems not necessarily conforming to Eq. (5), but we will consider a specific example of Hamiltonian engineering below where we seek to engineer $\hat{H}_{XXZ}$ into $\hat{H}_{ZZ}$.
@@
into an Ising model. Driving one of the spins in the XXZ interaction in the $xy$-plane, i.e., rotation about the $z$-axis, should cancel the XY-interaction term. To this end, consider a periodic driving pulse, $\hat{V}(t)$, of a single frequency given by $V_{0}\hat{\sigma}^{(3)}_{2}\cos(\omega t)$. We can use the Jacobi-Anger relation to show that it is possible drive $\hat{H}_{0}^{\text{S},2}$ into $\hat{H}_{ZZ}$ with $|V_{0}|\sim\omega$ (See, e.g., Ref. 38, for a related derivation). However, rigorous application of the Jacobi-Anger method is limited to a single frequency cosine driving pulse and application to general pulse profiles is not straightforward.
To engineer $\hat{H}_{0}^{\text{S},2}$ into an Ising model using a more general method (applicable to other pulse shapes), we can attempt to use the formalism of Ref. 23. This formalism can apply to pulse shapes of many frequencies, however the main results of Ref. 23 only apply when the pulse is weak, $|V_{0}|\ll\omega$. To see this we use we Ref. 23 by taking Fourier coefficients of $\hat{V}(t)$, and expand to lowest order in the effective Hamiltonian to arrive at a static effective model:
@@
we must require $\omega=|V_{0}|$, which violates the condition $|V_{0}|\ll\omega$. Similar issues arise in attempting pulses with other spatial profiels, e.g., experimentally relevant global pulses applied simultaneously to all sites. Using the main results of Ref. 23 thus requires a generalization to incorporate arbitrarily strong pulse profiles.
@@
As an additional example, consider the commonly applied global rectangular profiles. Fig. 1a shows an illustrative example. While the periodic driving is assumed to be fast with respect to other time scales set by $H_{0}$, the pulse itself introduces new times scales and other assumptions: i) Precision quantum simulation experiments using smooth approximations to rectangular pulses often need to account for the discrepancy between ideal and realized pulses–see, e.g., Ref. 13; ii) The pulse strength introduces an energy scale. The simplest approximations assume that the pulse strength is much smaller than $\omega$, but this assumption limits tunability of $\hat{H}_{\text{eff}}$ unless special assumptions are made; iii) The pulse width, $\tau_{w}$, must be assumed to be much shorter than $T$ but is bounded below by experimental constraints. For example, practical constraints on the $\tau_{w}\ll T$ assumption were studied in Ref. 9 using SMEs to show that finite pulse width led to an error in $\hat{H}_{\text{eff}}$ that scales as $\sim\tau_{w}/T$ with an impact on measurements of magnetization in their spin system to be as large as $10\%$.
These examples highlight the need for a rigorous general formalism with minimal assumptions on pulse shapes and controlled error. A step toward the generalization was considered in Appendix K of Ref. 23, where a single non-stroboscopic (global) pulse was considered with an arbitrary amplitude. Our central aim is to advance the approach discussed in Ref. 23 by constructing a formalism for Floquet engineering of non-stroboscopic effective Hamiltonians with a minimal set of assumptions on pulse shape and error controlled by only one small parameter, $\omega^{-1}$. We relax the assumption that the driving potential appears as a single pulse, and turn to the exposition of the formalism now.
@@ III Formalism
We implement an effective Hamiltonian formalism [23] by assuming a time-independent Hamiltonian, $\hat{H}_{0}$, with a periodic time-dependent perturbing potential, $\hat{V}(t)$. The formalism allows the transformation to a frame that yields a time-independent effective description enabled by expanding in powers of $\omega^{-1}$. The lowest order expansion yields effective Hamiltonians, $\hat{H}_{\text{\text{eff}}}$, and kick operators derived from a broad class of perturbing potentials without introducing additional time scales, imposing stroboscopic restrictions, or additional perturbative truncations. First, we will discuss the key assumptions defining the formalism. We then turn to the explicit theoretical description.
@@ III.1 Assumptions
    All operators are spanned by tensor products of generators of $SU(2)$, the usual Pauli spin operators.
@@
    The perturbing potential will represent global, time-ordered pulses which do not overlap in time.
@@ III.1.1 Cyclic, Periodic Perturbing Potential
As is common [39, 22, 24, 23], we define the perturbing potential to be periodic if there is a positive integer $N$ and duration $T$ such that the time-dependent potential $\hat{V}(t)$ satisfies:
@@ III.1.2 Operators as Linear Combinations of Pauli Matrices
thereby allowing us to separate $\hat{V}(t)$ into a sum of two terms: the term proportional to $f^{\alpha}_{j}(t)$ has an amplitude that is small compared to $\omega$ while the $\omega g^{\alpha}_{j}(t)$ term in general has an amplitude on the order of $\omega$. The case considered in Ref. 23 takes $g^{\alpha}_{j}(t)=0$ for all $t$ (with some special considerations in the appendix); we generalize this assumption in this work (Appendix A proves our generalization). Eq. (13) allows us to access a large class of pulse profiles.
Figure 1 shows example pulse shapes encompassed by Eq. (12). Panel (a) displays conventional pulse profiles [40] where pulse shapes are assumed to be rectangular for mathematical convenience in performing expansion integrals. The rectangular shape introduces a width time scale, $\tau_{w}$, implied by this type of pulse. Other time scales include the inverse pulse strength and those arising from derivatives near the edge of the rectangle (not shown). These time scales must be included in an accurate effective Hamiltonian formalism to mitigate errors. Furthermore, attempts to approximate ideal square pulse profiles in the laboratory can also lead to errors. Panel (b) shows another example class of pulse profiles where a discrete Fourier transform approximation to an otherwise continuous pulse profile is introduced. Here, again, the time scale $\tau_{w}$ is introduced along with other approximations that must be accounted for. Panel (c) depicts the general class of pulse profiles, Eq. (12), allowed by our formalism that include panels (a) and (b) as subsets. To accommodate pulses depicted in panel (c), our formalism must capture all assumptions regarding time scales and therefore allows pulses of arbitrary strength and duration provided $\omega^{-1}$ is kept as a small parameter.
III.1.3 Time-Ordering of Global Pulses
We also assume that pulses are time ordered, expressed as a restriction on $g^{\alpha}_{j}(t)$. In general, this restriction is not necessary, but it simplifies our formalism. For simplicity, we can consider the case where $f^{\alpha}_{j}(t)=0$ for all indices. We model the time-ordering by imposing a mathematical condition on the coefficients which ‘turns on’ one local spin operator at each site $j$ for each interval of $T$ so that we create subcycles of $[0,NT]$ of duration $T$. This can be expressed by requiring the $g^{\alpha}_{j}(t)$ to be piecewise continuous. Furthermore, we will also assume that the shape of the pulse remains the same as we cycle and also that the pulses are global. This means the functions have the form:

### v2

@@ I Introduction
Broadening the class of accessible quantum simulation models by improving the precision and applicability of engineering protocols will help experiments study models otherwise inaccessible to conventional numerical methods [BLOCH2012, BLATT2012, Georgescu2014, Gross2017, MONROE2021, MORGADO2021a, DALEY2022]. High frequency driving (Floquet engineering) with external fields has proven to be a useful experimental tool whereby otherwise static system Hamiltonians, $\hat{H}_{0}$, are perturbed with periodic external field pulses, $\hat{V}(t)$. The resulting driven Hamiltonian:
@@
and $T$ is a subcycle time defining periodicity such that $\hat{V}(t)$ is periodic in an integer multiple of $T$. $\omega^{-1}$ establishes a small expansion parameter compared to all other energy scales in $\hat{H}_{0}$. Interesting properties of eigenstates of $\hat{H}_{\text{eff}}$ can then be observed experimentally.
Recent experiments with Rydberg atoms [BLUVSTEIN2021a, GEIER2021c, scholl2022microwave, ZHAO2023], molecules [CHRISTAKIS2023b], and trapped ions [MORONG2023] have been able to dynamically engineer effective higher symmetry Heisenberg spin models from lower symmetry models. Ref. GEIER2021c used rectangular pulse profiles to engineer an effective $SU(2)$-symmetric Heisenberg interaction between a large number of Rydberg atoms initially lacking this symmetry in their interactions. Similarly, trapped ions were recently driven [MORONG2023] from an Ising limit into a higher-symmetry Heisenberg regime to replicate the Haldane-Shastry model [HALDANE1988b, SHASTRY1988]. Interest in engineering systems extends beyond transformations of underlying symmetry groups, or simulating $SU(2)$ Hamiltonians. For example, there have been proposals and experiments employing cold-atom systems to explore gauge theories and other field theories [halimeh2025cold, ott2021scalable, sundar2018synthetic, stannigel2014constrained], and trapped ions have been used to simulate complex open systems [barreiro2011open, MONROE2021].
The theory of Floquet engineering, and the associated high frequency expansions, has a long history (see Refs. [BLANES2009a, LESKES2010, goldman2014periodically, bukov2015universal] for reviews). There are two common methods to perform controlled expansions in inverse frequency that allow for effective engineering of broad classes of Hamiltonians. The first is the high frequency stroboscopic Magnus expansion (SME) [MAGNUS1954, bukov2015universal]. SMEs are commonly used in derivations of $\hat{H}_{\text{eff}}$, particularly in the context of average Hamiltonian theory [HAEBERLEN1968b, LESKES2010]. SMEs can be pushed to high order in inverse frequency to derive non-trivial effective models, e.g., multi-spin interactions [HERDMAN2010, choi2020robust], but often layer in additional assumptions. The second is the gauge invariant high frequency expansion (GIHFE) [rahav2003effective, goldman2014periodically, bukov2015universal, Abanin2017]. The GIHFE approach relies on the definition of a kick operator, $\hat{K}(t)$, from $\hat{V}(t)$ [goldman2014periodically]. The effective Hamiltonian and new gauge transformed state follow from:
@@
From these expressions we see that derivation of $\hat{H}_{\text{eff}}$ and derivation of the new states $\phi(t)$ from the original $\psi(t)$ then requires a systematic inverse frequency expansion of $\hat{K}(t)$ [rahav2003effective, goldman2014periodically, bukov2015universal] in addition to the expansion for $\hat{H}_{\text{eff}}$. Using $\hat{H}_{\text{eff}}$, $\hat{K}(t)$, and $\phi(t)$ one can then compute dressed (time averaged) observables [bukov2015universal].
GIHFEs, in contrast to SMEs, are often used to model sampling even at non-stroboscopic times by tracking the ensuing change in basis (gauge transformation) initiated by pulsing [Bukov2014]. Despite the distinction, one can recover average Hamiltonian theory from GIHFEs by restricting to stroboscopic sampling and applying the usual Magnus expansion [bukov2015universal]. The non-stroboscopic case can be constructed via a gauge transformation that depends on the initial sampling time. Assumptions underlying conventional GIHFE methods limit the space of available pulse profiles and therefore limit possible effective Hamiltonians.
We revisit the GIHFE formalism with the aim of broadening the set of available effective Hamiltonians in driven spin systems. The engineering method we construct generalizes the available pulse shapes considered in the existing literature [rahav2003effective, goldman2014periodically, bukov2015universal]. We consider periodic pulse structures such that $\omega^{-1}$ defines the only small parameter allowing otherwise arbitrary pulse shapes. Our formalism then allows for large pulse strengths, any pulse width that respects the high frequency approximation, and non-stroboscopic sampling. The pulse shape freedom we find arises because we are able to leverage the commutation relations of the Pauli spin matrices to perform controlled series expansions approximating $\hat{K}(t)$. Hamiltonian engineering then reduces to a functional optimized over the entire space of pulse shapes so that the desired effective Hamiltonian is realized.
To exemplify our approach, we apply our formalism to decrease the symmetry implied by native spin-spin interactions, motivated by the desire to construct (weighted) graph states [Hein2006a] in spin systems with native dipolar XXZ spin-spin interactions which characterize certain Rydberg atom [Saffman2010, NGUYEN2018] and molecule [Barnett2006, WALL2010a, muller2010, Wall2014] interactions. Since XXZ interactions do not commute in general, the protocol required to build weighted graph states introduces complexities that the Ising-type interaction avoids [Hein2006a, tanamoto2009efficient]. Requirements for building high precision weighted graph states with global pulses therefore motivates the specific example we use to demonstrate our formalism: engineering an XXZ interaction into an Ising interaction.
@@ II Spin Models
where $\hat{\sigma}^{(1)}_{j}$, $\hat{\sigma}^{(2)}_{j}$, and $\hat{\sigma}^{(3)}_{j}$ denote the usual $x$, $y$, and $z$ Pauli-spin matrices, respectively. $j$ and $j^{\prime}$ denote site (vertex) pairs in an arbitrary graph. $J^{1}_{jj^{\prime}}$, $J^{2}_{jj^{\prime}}$, and $J^{3}_{jj^{\prime}}$ denote spatially varying spin-spin interaction energies along the spin-$x$,$y$, and $z$ directions, respectively. Our formalism applies to more general models, but we will use $H_{0}$ as an example for applications. In specifying example applications, we restrict our attention to XXZ-type models with:
@@
Equation (4) encompasses well-known spin models. For $J_{z}=0$, Eq. (4) reduces to an XY-model, denoted $H_{XY}$. For $J_{\perp}=0$, we retrieve the Ising model, denoted $H_{ZZ}$. Eq. (5) implies that we may write $\hat{H}_{0}^{\text{S}}$ as $\hat{H}_{XXZ}=\hat{H}_{XY}+\hat{H}_{ZZ}$. Other choices of the interaction constants give rise to additional models, such as the Heisenberg model ($J_{\perp}=J_{z}$). Our formalism will hold for general spin systems not necessarily conforming to Eq. (5), but we will consider a specific example of Hamiltonian engineering below where we seek to engineer $\hat{H}_{XXZ}$ into $\hat{H}_{ZZ}$. Although we restrict to general 2-body interactions as given by Eq. (4), our considerations here can be generalized in a straightforward manner to higher order interactions.
@@
into an Ising model. Driving one of the spins in the XXZ interaction in the $xy$-plane, i.e., rotation about the $z$-axis, should cancel the XY-interaction term. To this end, consider a periodic driving pulse, $\hat{V}(t)$, of a single frequency given by $V_{0}\hat{\sigma}^{(3)}_{2}\cos(\omega t)$. We can use the Jacobi-Anger relation to show that it is possible drive $\hat{H}_{0}^{\text{S},2}$ into $\hat{H}_{ZZ}$ with $|V_{0}|\sim\omega$ (See, e.g., Ref. SOOMRO2022, for a related derivation). However, rigorous application of the Jacobi-Anger method is limited to a single frequency cosine driving pulse and application to general pulse profiles is not straightforward.
To engineer $\hat{H}_{0}^{\text{S},2}$ into an Ising model using a more general method (applicable to other pulse shapes), we can attempt to use the formalism of Ref. goldman2014periodically. This formalism can apply to pulse shapes of many frequencies, however the main results of Ref. goldman2014periodically only apply when the pulse is weak, $|V_{0}|\ll\omega$. To see this, we use Ref. goldman2014periodically by taking Fourier coefficients of $\hat{V}(t)$, and expand to lowest order in the effective Hamiltonian to arrive at a static effective model:
@@
we must require $\omega=|V_{0}|$, which violates the condition $|V_{0}|\ll\omega$. Similar issues arise in attempting pulses with other spatial profiles, e.g., experimentally relevant global pulses applied simultaneously to all sites. Using the main results of Ref. goldman2014periodically thus requires a generalization to incorporate arbitrarily strong pulse profiles.
@@
As an additional example, consider the commonly applied global rectangular profiles. Fig. 1a shows an illustrative example. While the periodic driving is assumed to be fast with respect to other time scales set by $\hat{H}_{0}$, the pulse itself introduces new times scales and other assumptions: i) Precision quantum simulation experiments using smooth approximations to rectangular pulses often need to account for the discrepancy between ideal and realized pulses–see, e.g., Ref. MORONG2023; ii) The pulse strength introduces an energy scale. The simplest approximations assume that the pulse strength is much smaller than $\omega$, but this assumption limits tunability of $\hat{H}_{\text{eff}}$ unless special assumptions are made; iii) The pulse width, $\tau_{w}$, must be assumed to be much shorter than $T$ but is bounded below by experimental constraints. For example, practical constraints on the $\tau_{w}\ll T$ assumption were studied in Ref. GEIER2021c using SMEs to show that finite pulse width led to an error in $\hat{H}_{\text{eff}}$ that scales as $\sim\tau_{w}/T$ with an impact on measurements of magnetization in their spin system to be as large as $10\%$.
These examples highlight the need for a rigorous general formalism with minimal assumptions on pulse shapes and controlled error. A step toward the generalization was considered in Appendix K of Ref. goldman2014periodically, where a single non-stroboscopic (global) pulse was considered with an arbitrary amplitude. Our central aim is to advance the approach discussed in Ref. goldman2014periodically by constructing a formalism for Floquet engineering of non-stroboscopic effective Hamiltonians with a minimal set of assumptions on pulse shape and error controlled by only one small parameter, $\omega^{-1}$. We relax the assumption that the driving potential appears as a single pulse, and turn to the exposition of the formalism now.
@@ III Formalism
We implement an effective Hamiltonian formalism [goldman2014periodically] by assuming a time-independent Hamiltonian, $\hat{H}_{0}$, with a periodic time-dependent perturbing potential, $\hat{V}(t)$. The formalism allows the transformation to a frame that yields a time-independent effective description enabled by expanding in powers of $\omega^{-1}$. The lowest order expansion yields effective Hamiltonians, $\hat{H}_{\text{\text{eff}}}$, and kick operators derived from a broad class of perturbing potentials without introducing additional time scales, imposing stroboscopic restrictions, or additional perturbative truncations. First, we will discuss the key assumptions defining the formalism. We then turn to the explicit theoretical description.
@@ III.1 Assumptions
    All observables are spanned by tensor products of generators of $SU(2)$, the usual Pauli spin operators, as well as the identity. In particular, the perturbing potential will assumed to be linear in the generators of $SU(2)$.
@@
    The lowest order kick operator will be assumed to be the time integral of the perturbing potential. As a special case, the perturbing potential will represent global, time-ordered pulses which do not overlap in time.
@@ III.1.1 Cyclic, Periodic Perturbing Potential
As is common [HAEBERLEN1968b, LESKES2010, bukov2015universal, goldman2014periodically], we define the perturbing potential to be periodic if there is a positive integer $N$ and duration $T$ such that the time-dependent potential $\hat{V}(t)$ satisfies:
@@ III.1.2 Operators as Linear Combinations of Pauli Matrices
thereby allowing us to separate $\hat{V}(t)$ into a sum of two terms: the term proportional to $f^{\alpha}_{j}(t)$ has an amplitude that is small compared to $\omega$ while the $\omega g^{\alpha}_{j}(t)$ term in general has an amplitude on the order of $\omega$. The case considered in Ref. goldman2014periodically takes $g^{\alpha}_{j}(t)=0$ for all $t$ (with some special considerations in the appendix); we generalize this assumption in this work (Appendix A proves our generalization). Eq. (13) allows us to access a large class of pulse profiles.
Figure 1 shows example pulse shapes encompassed by Eq. (12). Panel (a) displays conventional pulse profiles [Vandersypen2005] where pulse shapes are assumed to be rectangular for mathematical convenience in performing expansion integrals. The rectangular shape introduces a width time scale, $\tau_{w}$, implied by this type of pulse. Other time scales include the inverse pulse strength and those arising from derivatives near the edge of the rectangle (not shown). These time scales must be included in an accurate effective Hamiltonian formalism to mitigate errors. Furthermore, attempts to approximate ideal square pulse profiles in the laboratory can also lead to errors. Panel (b) shows another example class of pulse profiles where a discrete Fourier transform approximation to an otherwise continuous pulse profile is introduced. Here, again, the timescale $\tau_{w}$ is introduced along with other approximations that must be accounted for. Panel (c) depicts the general class of pulse profiles, Eq. (12), allowed by our formalism that include panels (a) and (b) as subsets. To accommodate pulses depicted in panel (c), our formalism must capture all assumptions regarding time scales and therefore allows pulses of arbitrary strength and duration provided $\omega^{-1}$ is kept as a small parameter.
III.1.3 Form of the Kick Operator
We assume that the relationship between the kick operator and the perturbing potential is given by
$\hat{K}^{(0)}(t)=\int^{t}\hat{V}(t^{\prime})dt^{\prime}=\sum_{\alpha,j}G^{\alpha}_{j}(t)\hat{\sigma}^{(\alpha)}_{j}$ (14)
with
$G^{\alpha}_{j}(t)=\omega\int^{t}g^{\alpha}_{j}(t^{\prime})dt^{\prime}.$ (15)
One can show that this form of the kick operator can be obtained by requiring that the terms of order $\mathcal{O}(\omega)$ sum to zero in the full expansion for $\hat{H}_{\text{eff}}$. [See Appendix B and Eq. (48).]
We make simplifying assumptions for the pulse. As a special case, we assume that pulses are time ordered, expressed as a restriction on $g^{\alpha}_{j}(t)$. In general, this restriction is not necessary, but it simplifies our formalism. For simplicity, we can consider the case where $f^{\alpha}_{j}(t)=0$ for all indices. We model the time-ordering by imposing a mathematical condition on the coefficients which ‘turns on’ one local spin operator at each site $j$ for each interval of $T$ so that we create subcycles of $[0,NT]$ of duration $T$. This can be expressed by requiring the $g^{\alpha}_{j}(t)$ to be piecewise continuous. Furthermore, we will also assume that the shape of the pulse remains the same as we cycle and also that the pulses are global. This means the functions have the form: